OpenStudy (anonymous):

Please help and explain this problem. Preferably step by step. Thanks!! A 3000kg train accelerates from rest to a velocity of 30m/s in 10s. The engine's power during this acceleration is ______.

OpenStudy (anonymous):

There no mention of friction. So, we can assume that the entire power generated by the engine is converted into its Kinetic Energy. Initial KE = 0J (zero initial velocity) Final KE = 1/2 * m * v^2 = 1/2 * 3000 * 30^2 = 1.35 * 10^6 J Therefore average power = (Final KE - Initial KE)/t = 1.35*10^6 / 10 = 135 kW